# Mitigating Attention Hacking in Preference-Based Reward Modeling via Interaction Distillation

### v1

@@ Abstract
The reward model (RM), as the core component of reinforcement learning from human feedback (RLHF) for large language models (LLMs), responsible for providing reward signals to generated responses. However, mainstream preference modeling in RM is inadequate in terms of token-level interaction, making its judgment signals vulnerable to being hacked by misallocated attention to context. This stems from two fundamental limitations: (1) Current preference modeling employs decoder-only architectures, where the unidirectional causal attention mechanism leads to forward-decaying intra-sequence attention within the prompt-response sequence. (2) The independent Siamese-encoding paradigm induces the absence of token-level inter-sequence attention between chosen and rejected sequences. To address this “attention hacking”, we propose “Interaction Distillation”, a novel training framework for more adequate preference modeling through attention-level optimization. The method introduces an interaction-based natural language understanding model as the teacher to provide sophisticated token interaction patterns via comprehensive attention, and guides the preference modeling to simulate teacher model’s interaction pattern through an attentional alignment objective. Through extensive experiments, interaction distillation has demonstrated its ability to provide more stable and generalizable reward signals compared to state-of-the-art RM optimization methods that target data noise, highlighting the attention hacking constitute a more fundamental limitation in RM.
Introduction
@@
Recently, reinforcement learning from human feedback (RLHF) has been widely adopted to enhance the alignment between large language model (LLM) behaviors and human preferences (Ouyang et al. 2022; Shen et al. 2023; Xu et al. 2025). The reward model (RM), as the core component of RLHF, is responsible for providing reward signals for LLM-generated responses (Wang et al. 2024a; Dou et al. 2024; Zhong et al. 2025), guiding LLMs to produce responses with higher reward scores. The current RM training approach first concatenates the prompt and response into a sequence, uses a LLM as the backbone to obtain sequence representations, and maps them to a scalar reward score via a linear layer. It then enforces the Bradley-Terry-based preference modeling (Bradley and Terry 1952) constraint to ensure that the chosen response receives a higher reward score than the rejected one (Schulman et al. 2017; Wang et al. 2024a). However, this preference modeling is inadequate at the token interaction level, as both intra-sequence and inter-sequence attention remain incomplete during training. This incompleteness makes the reward signals vulnerable to being hacked by misallocated attention to context.
@@
This “attention hacking” stems from two interaction limitations: (1) From intra-sequence attention perspective, existing preference-based RMs rely entirely on Transformer decoders as their backbone, where sequence attention is implemented via unidirectional causal multi-head attention. This forward-decaying attention mechanism prevents the RM from capturing interactions with preceding tokens in the sequence, resulting in disproportionate focus on the latter tokens of the prompt-response sequence. (2) From inter-sequence attention perspective, the independent Siamese-encoding paradigm in preference modeling leads to the absence of early interaction attention between chosen and rejected sequences. This design flaw prevents the RM from capturing token-level interactions, resulting in inadequate cross-sequence semantic perception.
@@
These token-level interaction limitations led us directly to consider the interaction-based natural language understanding (NLU) model, which is based on fine-tuned encoder-only Transformer (the BERT family (Devlin et al. 2019; He et al. 2022)) on NLU datasets (Camburu et al. 2018). It performs downstream classification tasks by jointly encoding the concatenation of two sequences. Since it processes the concatenated sequence, its bidirectional and global attention models the most comprehensive interactions between intra-sequence and inter-sequence tokens. Therefore, although interaction-based NLU models with the BERT family as their backbone are significantly smaller in scale compared to RMs based on LLMs, they serve as stronger teacher models in terms of attention.
@@
Inspired by this, we propose “Interaction Distillation”, a novel framework aiming to achieve more adequate token-level interaction in preference-based reward modeling, and the reward model trained with this method is termed Id-Rm. This framework consists of three steps: (1) Introducing an interaction-based NLU model as the teacher model, which unified encodes concatenated sequence pairs. Since it operates on concatenated sequences, its MHA can fully model both intra-sequence and inter-sequence token interactions. Consequently, we capture its global attention map as the learning objective for RM. (2) Leveraging the query and key matrices from preference modeling to simulate the teacher model’s comprehensive attention computation, guiding the standard RM to simulate the sophisticated interaction patterns of the teacher model. (3) Designing an attentional optimization objective to align the simulated attention maps from preference modeling with the real and comprehensive attention maps from the interaction-based NLU model, thereby bridging the gap in token-level interaction between RM and the teacher model.
@@
The current state-of-the-art (SOTA) preference-based reward modeling optimization methods primarily address data noise (Touvron et al. 2023; Wang et al. 2024b; Rame et al. 2024; Miao et al. 2024). In contrast, Id-Rm delivers more stable and generalizable reward signals without introducing inference costs, with its advantages manifested in two key aspects: (1) For RLHF tasks (Völske et al. 2017; Bai et al. 2022), the Id-Rm-optimized policy models generate responses that best align with human preferences. (2) In out-of-distribution preference perception tasks (Lambert et al. 2024; Zhou et al. 2024), Id-Rm achieves the highest average accuracy. These results highlight that the intrinsic deficiencies in RM’s attention mechanisms constitute a more fundamental limiting factor than data noise. In conclusion, our main contributions are summarized as follows:
@@
Attention Hacking in Reward Modeling
In this section, we first introduce the training objective for preference-based reward models. We then analyze how both intra- and inter-sequence attention in preference modeling are directly linked to the reward model’s decision. However, the inherent nature of preference modeling is inadequate at the token interaction level, making reward signals vulnerable to being hacked by misallocated attention to context.
@@
Preference-based Reward Modeling
[…]
@@
Perspective from Intra-Sequence Attention
@@
Perspective from Inter-Sequence Attention
@@
Method
In this section, we propose “Interaction Distillation” to address the attention hacking and achieve more adequate preference modeling in reward model, and the reward model trained with this method is termed Id-Rm. As shown in Figure 3, interaction distillation consists of three steps: (1) introducing an interaction-based natural language understanding (NLU) model as a teacher to extract comprehensive attention maps across prompt-response sequences; (2) performing identical attention computation in preference modeling as in the teacher model to simulate the teacher’s interaction patterns; (3) introducing an attentional optimization objective $\mathcal{L}_{\text{ID}}$ to align the comprehensive attention maps from the NLU model with the simulated attention maps from preference modeling, as formulated in Equation 3.
@@
Interaction based NLU Model as the Teacher
As shown in Figure 3, the interaction-based NLI model, composed of an encoder-only Transformer and a linear head, has emerged as the unified paradigm for natural language understanding tasks (Devlin et al. 2019), with the BERT family being the most prominent representative (Devlin et al. 2019; He et al. 2022; Zang and Liu 2024a). It jointly encodes concatenated input sequences, pools representations tokens, and adapts them to classification (e.g. {entailment, neutral, contradiction}) via a linear layer. In our framework, we concatenate the chosen and rejected prompt-response pairs into a sequence $[s_{c};s_{r}]$, which is mapped representation $\mathbf{H}=[\mathbf{H}_{c};\mathbf{H}_{r}]\in\mathbb{R}^{(l_{c}+l_{r})\times d}$. $\mathbf{H}$ is enhanced through the weighted fusion of multi-block attention maps $\mathbf{A}$ (solid line map) in NLI model, where $\mathbf{A}$ is formulated in Equation 4.
@@
$=\text{softmax}(\begin{bmatrix}\mathbf{Q}_{c}\mathbf{K}^{\top}_{c}/\sqrt{d}&\mathbf{Q}_{c}\mathbf{K}^{\top}_{r}/\sqrt{d}\\ \mathbf{Q}_{r}\mathbf{K}^{\top}_{c}/\sqrt{d}&\mathbf{Q}_{r},\mathbf{K}_{r}/\sqrt{d}\end{bmatrix}$
=$\overset{\rm def}{=}\begin{bmatrix}\mathbf{A}_{c\rightarrow c}&\mathbf{A}_{c\rightarrow r}\\ \mathbf{A}_{r\rightarrow c}&\mathbf{A}_{r\rightarrow r}\end{bmatrix}$
@@
Simulated Interaction in Preference Modeling
@@
Helpful Harmless Chat Chat-Hard Safety Reasoning
DPO 57.87 63.44 92.84 49.67 73.87 57.67 65.89
BT-RM 56.63 66.71 90.03 49.99 74.21 61.42 66.50
AM-RM 57.27 65.13 91.24 50.92 74.43 62.75 66.96
LF-RM 57.78 66.38 92.45 51.21 75.99 62.19 67.67
WARM 56.84 67.58 91.02 52.98 77.42 63.44 68.50
InfoRM 58.90 66.02 91.45 53.12 79.12 62.90 68.59
Id-Rm 59.07 68.11 92.23 54.56 78.34 67.23 69.92
@@
Helpful Harmless $\Delta$ (Avg.) Helpful Harmless $\Delta$ (Avg.)
Teacher: DeBERTa w/ tuning on SNLI (base) 60.66 65.63 - 59.07 68.11 -
Teacher: DeBERTa w/o tuning on SNLI Ablation 54.14 61.33 -5.41 56.24 66.83 -2.06
Teacher: BERT w/ tuning on SNLI 60.74 65.23 -0.16 58.20 68.77 -0.11
Teacher: BERT w/o tuning on SNLI 54.92 62.09 -4.64 56.55 67.21 -1.71
Loss: w/o Intra-sequence distillation 52.45 57.23 -8.30 57.04 67.19 -1.48
Loss: w/o Inter-sequence distillation 58.44 62.98 -2.44 58.22 67.86 -0.55
w/ Adaptive margining Compatibility 61.89 66.79 1.20 61.22 70.03 2.04
w/ Label flipping 62.07 66.21 1.00 62.19 69.44 2.22
w/ Weight averaging 63.03 66.52 1.63 62.01 70.02 2.43
Table 3: Ablation and compatibility study results of Id-Rm. We report the win rate (%) of Id-Rm against BT-RM on HH-RLHF and the OOD accuracy (%) on RMB. $\Delta$ (Avg.) indicates the average change relative to the base configuration.
Experiments
Experimental Setup
In this study, we employed LLaMA3 (Grattafiori et al. 2024) with 8 billion parameters as the backbone for the preference-based reward model, and adopted an interaction-based NLU model based on DeBERTa-large (He et al. 2022) as the teacher model. The teacher model was fine-tuned on natural language understanding dataset SNLI (Bowman et al. 2015). To evaluate stability and generalizability of the reward signals provided by Id-Rm, we conducted a series of experiments in both reinforcement learning from human feedback (RLHF) and out-of-distribution (OOD) preference perception. The between-group differences in the experimental results reported in this study were all statistically significant at the 5% confidence level, as verified by Wilcoxon signed-rank tests.
RLHF Setting.
In RLHF tasks, we use Id-Rm to provide supervision signals for the policy model and optimize the policy model using Proximal Policy Optimization (PPO) (Schulman et al. 2017). The policy model is initialized via supervised fine-tuning (SFT) beforehand. For the general dialogue task, we used the HH-RLHF dataset (Bai et al. 2022) for PPO optimization, focusing on the helpful and harmless perspectives. For the summarization task, we employed the TL;DR dataset (Völske et al. 2017).
For evaluation, we used the policy model trained with Id-Rm to answer questions from the test set, while the SFT model and the policy model trained via PPO with Bradley-Terry RM served as opponents answering the same test questions. Existing research demonstrates strong correlation between GPT-4o’s response assessments and human evaluation results (Zheng et al. 2023; Chang et al. 2024). Thus, we evaluated their responses using GPT-4o (Achiam et al. 2023) and calculated the win rate of Id-Rm, where the win rate is calculated as the proportion of winning samples to the total number of winning and losing samples.
OOD Preference Perception Setting.
In OOD preference perception, to explore whether Id-Rm can provide generalizable signals in out-of-distribution tasks, we performed inference-time verification on the world reward model benchmark RMB (Zhou et al. 2024) and Reward Bench (Lambert et al. 2024), using accuracy for evaluation. Among them, RMB includes two dimensions: helpfulness and harmlessness, while Reward Bench comprises four subtasks: Chat, Chat-Hard, Safety, and Reasoning.
Baselines.
We employ SFT model and Direct Preference Optimization (DPO) model (Rafailov et al. 2023) as baselines for RLHF tasks. Additionally, we select SOTA reward model optimization methods as baselines: Bradley-Terry-based BT-RM (Schulman et al. 2017), adaptive-margin based AM-RM (Touvron et al. 2023), label-flipping-based LF-RM (Burns et al. 2024; Wang et al. 2024b), weight averaging based WARM (Rame et al. 2024), and information-bottleneck-based InfoRM (Miao et al. 2024), which primarily address noise handling in the data.
Main Results
RLHF tasks.

### v3

# Alleviating Attention Hacking in Discriminative Reward Modeling through Interaction Distillation

@@ Abstract
The reward model (RM), as the core component of reinforcement learning from human feedback (RLHF) for large language models (LLMs), responsible for providing reward signals to generated responses. However, the mainstream discriminative reward modeling is inadequate in terms of token-level interaction, making its judgment signals vulnerable to being hacked by misallocated attention to context. This stems from two fundamental limitations: (1) Current preference modeling employs decoder-only architectures, where the unidirectional causal attention mechanism leads to forward-decaying intra-sequence attention within the prompt-response sequence. (2) The independent Siamese-encoding paradigm induces the absence of token-level inter-sequence attention between chosen and rejected sequences. To address this “attention hacking”, we propose “Interaction Distillation”, a novel training framework for more adequate discriminative reward modeling via attention-level optimization. The method introduces an interaction-based natural language understanding model as the teacher to provide sophisticated token interaction patterns via comprehensive attention, and guides the reward modeling to simulate teacher model’s interaction pattern through an attentional alignment objective. Through extensive experiments, interaction distillation has demonstrated its ability to provide more stable and generalizable reward signals compared to state-of-the-art RM optimization methods that target data noise, highlighting the attention hacking constitute a more fundamental limitation in discriminative RM.
1 Introduction
@@
Recently, reinforcement learning from human feedback (RLHF) has been widely adopted to enhance the alignment between large language model (LLM) behaviors and human preferences Ouyang et al. (2022); Shen et al. (2023); Xu et al. (2025). The reward model (RM), as the core component of RLHF, is responsible for providing reward signals for LLM-generated responses Wang et al. (2024a); Dou et al. (2024); Zhong et al. (2025), guiding LLMs to produce responses with higher reward scores. The mainstream discriminative RM training first concatenates the prompt and response into a sequence, uses a LLM as the backbone to obtain sequence representations, and maps them to a scalar reward score via a linear layer. It then enforces the Bradley-Terry-based preference modeling Bradley and Terry (1952) constraint to ensure that the chosen response receives a higher reward score than the rejected one Schulman et al. (2017); Wang et al. (2024a). However, this preference modeling is inadequate at the token interaction level, as both intra-sequence and inter-sequence attention remain incomplete during training. This incompleteness makes the reward signals vulnerable to being hacked by misallocated attention to context.
@@
This “attention hacking” stems from two interaction limitations: (1) From intra-sequence attention perspective, existing discriminative RMs rely entirely on Transformer decoders as their backbone, where sequence attention is implemented via unidirectional causal multi-head attention. This forward-decaying attention mechanism prevents the RM from capturing interactions with preceding tokens in the sequence, resulting in disproportionate focus on the latter tokens of the prompt-response sequence. (2) From inter-sequence attention perspective, the independent Siamese-encoding paradigm in preference modeling leads to the absence of early interaction attention between chosen and rejected sequences. This design flaw prevents the RM from capturing token-level interactions, resulting in inadequate cross-sequence semantic perception.
@@
These token-level interaction limitations led us directly to consider the interaction-based natural language understanding (NLU) model, which is based on fine-tuned encoder-only Transformer (the BERT family Devlin et al. (2019); He et al. (2022)) on NLU datasets Camburu et al. (2018). It performs downstream classification tasks by jointly encoding the concatenation of two sequences. Since it processes the concatenated sequence, its bidirectional and global attention models the most comprehensive interactions between intra-sequence and inter-sequence tokens. Therefore, although interaction-based NLU models with the BERT family as their backbone are significantly smaller in scale compared to RMs based on LLMs, they serve as stronger teacher models in terms of attention.
@@
Inspired by this and knowledge distillation Li et al. (2022); Lu et al. (2022), we propose “Interaction Distillation”, a novel framework aiming to achieve more adequate token-level interaction in discriminative reward modeling, and the reward model trained with this method is termed Id-Rm. This framework consists of three steps: (1) Introducing an interaction-based NLU model as the teacher model, which unified encodes concatenated sequence pairs. Since it operates on concatenated sequences, its MHA can fully model both intra-sequence and inter-sequence token interactions. Consequently, we capture its global attention map as the learning objective for RM. (2) Leveraging the query and key matrices from preference modeling to simulate the teacher model’s comprehensive attention computation, guiding the standard RM to simulate the sophisticated interaction patterns of the teacher model. (3) Designing an attentional optimization objective to align the simulated attention maps from preference modeling with the real and comprehensive attention maps from the interaction-based NLU model, thereby bridging the gap in token-level interaction between RM and the teacher model.
@@
The current state-of-the-art (SOTA) discriminative reward modeling optimization methods primarily address data noise Touvron et al. (2023); Wang et al. (2024b); Rame et al. (2024); Miao et al. (2024). In contrast, Id-Rm delivers more stable and generalizable reward signals without introducing inference costs, with its advantages manifested in two key aspects: (1) For RLHF tasks Völske et al. (2017); Bai et al. (2022), the Id-Rm-optimized policy models generate responses that best align with human preferences. (2) In out-of-distribution preference perception tasks Lambert et al. (2024); Zhou et al. (2024), Id-Rm achieves the highest average accuracy. These results highlight that the intrinsic deficiencies in RM’s attention mechanisms constitute a more fundamental limiting factor than data noise. In conclusion, our main contributions are summarized as follows:
@@
2 Attention Hacking in Discriminative RM
In this section, we first introduce the training objective for discriminative reward models. We then analyze how both intra- and inter-sequence attention in preference modeling are directly linked to the reward model’s decision. However, the inherent nature of preference modeling is inadequate at the token interaction level, making reward signals vulnerable to being hacked by misallocated attention to context.
@@ 2 Attention Hacking in Discriminative RM
2.1 Discriminative Reward Modeling
Generally, the $\theta$ parameterized discriminative reward model $\mathcal{R}_{\theta}$ consists of a decoder-only Transformer and a linear head, which takes the prompt $x$ and response $y$ as input. Formally, let the Transformer’s input sequence be $s=[\langle\text{bos}\rangle;x;\langle\text{eos}\rangle;y;\langle\text{eos}\rangle]$, where $\langle\text{bos}\rangle$ and $\langle\text{eos}\rangle$ denote special tokens. Its high-dimensional representation is denoted as $\mathbf{H}^{\text{PM}}=[\mathbf{h}_{1},\mathbf{h_{2}},...,\mathbf{h}_{l}]\in\mathbb{R}^{l\times d}$, where $l$ is the input sequence length and $d$ is the representation dimension. Typically, the last token’s representation is projected through a linear layer to obtain the reward score $\mathcal{R}_{\theta}(x,y)$, formulated in Equation 1.
@@ 2.1 Discriminative Reward Modeling
2.2 Perspective from Intra-Sequence Attention
@@
2.3 Perspective from Inter-Sequence Attention
@@
3 Method
Inspired by attention-level knowledge distillation Li et al. (2022); Lu et al. (2022), we propose “Interaction Distillation” to address the attention hacking and achieve more adequate preference modeling in discriminative RM, and the reward model trained with this method is termed Id-Rm. As shown in Figure 3, interaction distillation consists of three steps: (1) introducing an interaction-based natural language understanding (NLU) model as a teacher to extract comprehensive attention maps across prompt-response sequences; (2) performing identical attention computation in preference modeling as in the teacher model to simulate the teacher’s interaction patterns; (3) introducing an attentional optimization objective $\mathcal{L}_{\text{ID}}$ to align the comprehensive attention maps from the NLU model with the simulated attention maps from preference modeling, as formulated in Equation 3.
@@ 3 Method
3.1 Interaction based NLU Model as the Teacher
As shown in Figure 3, the interaction-based NLI model, composed of an encoder-only Transformer and a linear head, has emerged as the unified paradigm for natural language understanding tasks Devlin et al. (2019), with the BERT family being the most prominent representative Devlin et al. (2019); He et al. (2022). It jointly encodes concatenated input sequences, pools representations tokens, and adapts them to classification (e.g. {entailment, neutral, contradiction}) via a linear layer. In our framework, we concatenate the chosen and rejected prompt-response pairs into a sequence $s_{c};s_{r}]$, which is mapped representation $\mathbf{H}=[\mathbf{H}_{c};\mathbf{H}_{r}]\in\mathbb{R}^{(l_{c}+l_{r})\times d}$. $\mathbf{H}$ is enhanced through the weighted fusion of multi-block attention maps $\mathbf{A}$ (solid line map in Figure 4) in NLI model, where $\mathbf{A}$ is formulated in Equation 4.
@@ 3.1 Interaction based NLU Model as the Teacher
$=\text{softmax}(\begin{bmatrix}\mathbf{Q}_{c}\mathbf{K}^{\top}_{c}/\sqrt{d}&\mathbf{Q}_{c}\mathbf{K}^{\top}_{r}/\sqrt{d}\\ \mathbf{Q}_{r}\mathbf{K}^{\top}_{c}/\sqrt{d}&\mathbf{Q}_{r}\mathbf{K}_{r}/\sqrt{d}\end{bmatrix}$
$\overset{\rm def}{=}\begin{bmatrix}\mathbf{A}_{c\rightarrow c}&\mathbf{A}_{c\rightarrow r}\\ \mathbf{A}_{r\rightarrow c}&\mathbf{A}_{r\rightarrow r}\end{bmatrix}$
@@
Figure 4: Attention transfer mechanism in interaction distillation
3.2 Simulated Interaction in Preference Modeling
@@ 3.3 Interaction Distillation
Harmless Helpful Chat Chat-Hard Safety Reasoning
DPO 57.87 53.44 92.84 49.67 73.87 57.67 64.22
BT-RM 56.63 56.71 90.03 49.99 74.21 61.42 64.83
AM-RM 57.27 55.13 91.24 50.92 74.43 62.75 65.29
LF-RM 57.78 56.38 92.45 51.21 75.99 62.19 66.00
WARM 56.84 57.58 91.02 52.98 77.42 63.44 66.83
InfoRM 58.90 56.02 91.45 53.12 79.12 62.90 66.92
Id-Rm 59.07 57.11 92.23 54.56 78.34 67.23 68.08
@@
4 Experiments
4.1 Experimental Setup
In this study, we employed LLaMA3 Grattafiori et al. (2024) with 8 billion parameters as the backbone for the di s reward model, and adopted an interaction-based NLU model based on DeBERTa-large He et al. (2022) as the teacher model. The teacher model was fine-tuned on natural language understanding dataset SNLI Bowman et al. (2015). To evaluate stability and generalizability of the reward signals provided by Id-Rm, we conducted a series of experiments in both reinforcement learning from human feedback (RLHF) and out-of-distribution (OOD) preference perception. The between-group differences in the experimental results reported in this study were all statistically significant at the 5% confidence level, as verified by Wilcoxon signed-rank tests [^1] Detailed setup, implementations, and case studies provided in the supplementary materials.
4.1.1 RLHF Setting.
In RLHF tasks, we use Id-Rm to provide supervision signals for the policy model and optimize the policy model using Proximal Policy Optimization (PPO) Schulman et al. (2017). The policy model is initialized via supervised fine-tuning (SFT) beforehand. For the general dialogue task, we used the HH-RLHF dataset Bai et al. (2022) for PPO optimization, focusing on the helpful and harmless perspectives. For the summarization task, we employed the TL;DR dataset Völske et al. (2017).
For evaluation, we used the policy model trained with Id-Rm to answer questions from the test set, while the SFT model and the policy model trained via PPO with Bradley-Terry RM served as opponents answering the same test questions. Existing research demonstrates strong correlation between GPT-4o’s response assessments and human evaluation results Zheng et al. (2023); Chang et al. (2024). Thus, we evaluated their responses using GPT-4o Achiam et al. (2023) and calculated the win rate of Id-Rm, where the win rate is calculated as the proportion of winning samples to the total number of winning and losing samples.
4.1.2 OOD Preference Perception Setting.
In OOD preference perception, to explore whether Id-Rm can provide generalizable signals in out-of-distribution tasks, we performed inference-time verification on the world reward model benchmark RMB Zhou et al. (2024) and Reward Bench Lambert et al. (2024), using accuracy for evaluation. Among them, RMB includes two dimensions: helpfulness and harmlessness, while Reward Bench comprises four subtasks: Chat, Chat-Hard, Safety, and Reasoning.
4.1.3 Baselines.
We employ SFT model and Direct Preference Optimization (DPO) model Rafailov et al. (2023) as baselines for RLHF tasks. Additionally, we select SOTA reward model optimization methods as baselines: Bradley-Terry-based BT-RM Schulman et al. (2017), adaptive-margin based AM-RM Touvron et al. (2023), label-flipping-based LF-RM Burns et al. (2024); Wang et al. (2024b), weight averaging based WARM Rame et al. (2024), and information-bottleneck-based InfoRM Miao et al. (2024), which primarily address noise handling in the data.
@@ 4.1 Experimental Setup
Helpful Harmless $\Delta$ (Avg.) Harmless Helpful $\Delta$ (Avg.)
Teacher: DeBERTa w/ tuning on SNLI (base) 60.66 65.63 - 59.07 57.11 -
Teacher: DeBERTa w/o tuning on SNLI Ablation 54.14 61.33 -5.41 56.24 56.83 -2.06
Teacher: BERT w/ tuning on SNLI 60.74 65.23 -0.16 58.20 58.77 -0.11
Teacher: BERT w/o tuning on SNLI 54.92 62.09 -4.64 56.55 57.21 -1.71
Loss: w/o Intra-sequence distillation 52.45 57.23 -8.30 57.04 57.19 -1.48
Loss: w/o Inter-sequence distillation 58.44 62.98 -2.44 58.22 57.86 -0.55
w/ Adaptive margining Compatibility 61.89 66.79 1.20 61.22 60.03 2.04
w/ Label flipping 62.07 66.21 1.00 62.19 59.44 2.22
w/ Weight averaging 63.03 66.52 1.63 62.01 60.02 2.43
Table 3: Ablation and compatibility study results of Id-Rm. We report the win rate (%) of Id-Rm against BT-RM on $H$-RLHF and the OOD accuracy (%) on RMB. $\Delta$ (Avg.) indicates the average change relative to the base configuration.
4.2 Main Results
4.2.1 RLHF tasks